OpenStudy (anonymous):

The odds against Lounging Larry winning the horse race are 12:5. Determine the probability that Lounging Larry wins the horse race. 5/12 12/5 12/17 5/17 17/5

OpenStudy (anonymous):

i think it is 5/12 but that is larrys odds of winning

OpenStudy (anonymous):

because if the horse odds 12 to 5 then larrys odds is 5 in12

OpenStudy (anonymous):

hope i am right

OpenStudy (kropot72):

The odds are as follows: P(Larry does not win) : P(larry wins) = 12 : 5 Therefore P(Larry does not win) = 12/17 P(Larry wins) = 5/17